# BOTH2Hands: Inferring 3D Hands from Both Text Prompts and Body Dynamics

### 20 Dec 2023 v3

@@ A.3 Data Process
$0.9/{0.1}$ $0.032^{\pm 0.022}$ $0.063^{\pm 0.028}$ $0.093^{\pm 0.034}$ ${\color[rgb]{1,0,0}0.198}^{\pm 0.022}$ $1.402^{\pm 0.012}$ $3.968^{\pm 0.074}$ $1.298^{\pm 0.124}$
$0.8/{0.2}$ $0.037^{\pm 0.014}$ ${\color[rgb]{0,0,1}0.075}^{\pm 0.020}$ ${\color[rgb]{1,0,0}0.115}^{\pm 0.028}$ $0.201^{\pm 0.020}$ ${\color[rgb]{0,0,1}1.392}^{\pm 0.008}$ $3.969^{\pm 0.082}$ ${\color[rgb]{1,0,0}1.312}^{\pm 0.034}$
$0.7/{0.3}$ $0.031^{\pm 0.020}$ $0.060^{\pm 0.024}$ $0.088^{\pm 0.015}$ ${\color[rgb]{0,0,1}0.199}^{\pm 0.020}$ $1.404^{\pm 0.008}$ $3.968^{\pm 0.060}$ ${\color[rgb]{0,0,1}1.303}^{\pm 0.104}$
@@
$0.4/{0.6}$ ${\color[rgb]{0,0,1}0.038}^{\pm 0.010}$ $0.074^{\pm 0.016}$ $0.104^{\pm 0.024}$ $0.237^{\pm 0.038}$ ${\color[rgb]{1,0,0}1.391}^{\pm 0.010}$ $3.939^{\pm 0.070}$ $1.250^{\pm 0.136}$
@@
$0.1/{0.9}$ ${\color[rgb]{1,0,0}0.041}^{\pm 0.020}$ ${\color[rgb]{1,0,0}0.079}^{\pm 0.020}$ ${\color[rgb]{0,0,1}0.113}^{\pm 0.032}$ $0.233^{\pm 0.026}$ $1.392^{\pm 0.010}$ $3.934^{\pm 0.052}$ $1.223^{\pm 0.268}$
@@ B.1 Hyperparameter experiments
We established 11 sets of hyperparameter experiments with an interval of 0.1, starting at 0, under the premise that $w_{B}+w_{T}=1$. We limit $w_{B}+w_{T}=1$ to ensure the cross-attention transformer layer returns the same order of magnitude of loss as the previous diffusion process. Tab. 5 shows detailed results under different hyperparameter settings. The evaluation metrics are composed of the following sections: Motion-retrieval precision (R Precision), where we adhere to the top 3 results, is used to measure the alignment degree between conditions and motions. Fréchet Inception Distance (FID) [17] is employed to evaluate the feature distribution between the generated actions and the ground truth motion. Multi-modal Distance (MM-Dist) calculates the distance between hand motions and body text conditions. Diversity assesses the richness of generation motion by calculating the variance of data. Multimodality (MModality) measures the variety of generated motions under the same input conditions. The experimental results demonstrated that utilizing solely the body or text as the condition, equal to setting $w_{T}=0$ or $w_{B}=0$ in the experiment, can not get the best results. As the weight $w_{T}$ of the text condition increases, the alignment performance (R Precision) between motion and condition improves, but the authenticity of the motion (FID) deteriorates. As the weight $w_{B}$ of the body condition increases, the authenticity of the motion (FID) improves, but the Multi-modal Distance (MM-Dist) also increases, indicating that the distance between the generated hand motion result and the condition has grown farther in latent space. Under the experiment results, we decided to use a group with a good match degree between motion and condition, and relatively realistic motion, specifically setting the hyperparameters of the pipeline as $w_{B}=0.8$ and $w_{T}=0.2$.
@@ B.2 More evaluation results
We present more visualization results of comparing methods to further validate the effectiveness of the BOTH2Hands algorithm. As shown in Fig. 11, under the same text condition but different body conditions, our method can still generate hand results that correspond to the text and body. We have also illustrated more cross-validation visualization results. As depicted in Fig. 12, under the challenge of more vague text descriptions and professional dance movements, our dataset can still provide lively hand gestures that align with text controls, demonstrating the broad generalization capabilities of the BOTH57M dataset. Finally, we present a gallery of our inference results for the body text to hand task in Fig. 13.
@@ C Applications
In case the BOTH57M dataset encompasses rich annotations of both hand and body movements, our BOTH2Hands algorithm, trained on this dataset, has demonstrated excellent generalization capabilities even on other human body datasets that only offer body-level descriptions. We experimented with feeding body and text inputs from the HumanML3D [13] dataset into BOTH2Hands and succeeded in generating hand movements that are realistically vivid. In a bid to challenge the generation of hand movements in more professional actions, we proceeded to perform inference on InterHuman [25], a dataset featuring professional movements like ballet, boxing, etc. Fig. 14 illustrates how our method can effectively augment data for existing body and text datasets.
@@ D Limitation and Broader Impact
We discuss the border impact and limitations in this section.
Broader Impact. We are the first to propose generating fine-grained two-hand motions through implicit control from real-world body input and user-defined explicit text control. This provides a direction that can be studied for user-customized motion generation. Furthermore, the pre-trained BOTH2Hands, based on the hand synthesis through the parallel diffusion structure and large-scale motion data training on the BOTH57M, can enhance data for large body motion datasets with textual descriptions like HumanML3D and InterHuman. It also contributes to balancing control signals during the multi-signal control diffusion process. Furthermore, text-controllable and body-aligned hand synthesis has potential in many practical application scenarios, such as Virtual Reality (VR), Augmented Reality (AR) games, animations, and human communication studies.
Limitation. Despite the fact that this work has made significant progress on the novel task of generating hand motions that align with body movements and provide an intuitive, explicit control method, the method still has some limitations. Firstly, the fine-grained spatial control provided by text descriptions is sufficient, while there is a lack of temporal alignment methods. Secondly, we focused on the connection between hand actions and the body as a whole but did not take into account how different parts of the body might separately affect the hands. Future work could possibly focus on aligning text control temporally and exploring how individual body parts relate to hand movements.

### 09 Apr 2024 v4

@@ A.6 Data Statics
$0.9/{0.1}$ $0.032^{\pm 0.022}$ $0.063^{\pm 0.028}$ $0.093^{\pm 0.034}$ ${\color[rgb]{1,0,0}\textbf{0.198}}^{\pm 0.022}$ $1.402^{\pm 0.012}$ $3.968^{\pm 0.074}$ $1.298^{\pm 0.124}$
$0.8/{0.2}$ $0.037^{\pm 0.014}$ ${\color[rgb]{0,0,1}\textbf{0.075}}^{\pm 0.020}$ ${\color[rgb]{1,0,0}\textbf{0.115}}^{\pm 0.028}$ $0.201^{\pm 0.020}$ ${\color[rgb]{0,0,1}\textbf{1.392}}^{\pm 0.008}$ $3.969^{\pm 0.082}$ ${\color[rgb]{1,0,0}\textbf{1.312}}^{\pm 0.034}$
$0.7/{0.3}$ $0.031^{\pm 0.020}$ $0.060^{\pm 0.024}$ $0.088^{\pm 0.015}$ ${\color[rgb]{0,0,1}\textbf{0.199}}^{\pm 0.020}$ $1.404^{\pm 0.008}$ $3.968^{\pm 0.060}$ ${\color[rgb]{0,0,1}\textbf{1.303}}^{\pm 0.104}$
@@
$0.4/{0.6}$ ${\color[rgb]{0,0,1}\textbf{0.038}}^{\pm 0.010}$ $0.074^{\pm 0.016}$ $0.104^{\pm 0.024}$ $0.237^{\pm 0.038}$ ${\color[rgb]{1,0,0}\textbf{1.391}}^{\pm 0.010}$ $3.939^{\pm 0.070}$ $1.250^{\pm 0.136}$
@@
$0.1/{0.9}$ ${\color[rgb]{1,0,0}\textbf{0.041}}^{\pm 0.020}$ ${\color[rgb]{1,0,0}\textbf{0.079}}^{\pm 0.020}$ ${\color[rgb]{0,0,1}\textbf{0.113}}^{\pm 0.032}$ $0.233^{\pm 0.026}$ $1.392^{\pm 0.010}$ $3.934^{\pm 0.052}$ $1.223^{\pm 0.268}$
@@ B.1 Hyperparameter experiments
We establish 11 sets of hyperparameter experiments with an interval of 0.1, starting at 0, under the premise that $w_{B}+w_{T}=1$. We limit $w_{B}+w_{T}=1$ to ensure the cross-attention transformer layer returns the same order of magnitude of loss as the previous diffusion process. Tab. 6 shows detailed results under different hyperparameter settings. The evaluation metrics are composed of the following sections: Motion-retrieval precision (R Precision), which we adhere to the top 3 results, is used to measure the alignment degree between conditions and motions. Fréchet Inception Distance (FID) [19] is employed to evaluate the feature distribution between the generated actions and the ground truth motion. Multi-modal Distance (MM-Dist) calculates the distance between hand motions and body text conditions. Diversity assesses the richness of generation motion by calculating the variance of data. Multimodality (MModality) measures the variety of generated motions under the same input conditions. The experimental results demonstrate that utilizing solely the body or text as the condition, corresponding to setting $w_{T}=0$ or $w_{B}=0$ in the experiment, can not get the best results. As the weight of the text condition $w_{T}$ increases, the alignment performance (R Precision) between motion and condition improves, but the authenticity of the motion (FID) deteriorates. As the weight of the body condition $w_{B}$ increases, the authenticity of the motion (FID) improves, but the Multi-modal Distance (MM-Dist) also increases, indicating that the distance between the generated hand motion result and the condition has grown farther in latent space. Under the experiment results, we decide to use a group with a good match degree between motion and condition, and relatively realistic motion, specifically setting the hyperparameters of the pipeline as $w_{B}=0.8$ and $w_{T}=0.2$.
@@ B.2 More evaluation results
We present more visualization results of comparing methods to further validate the effectiveness of the BOTH2Hands algorithm. As shown in Fig. 11, under the same text condition but different body conditions, our method can still generate hand results that correspond to the text and body. We also illustrate more cross-validation visualization results. As depicted in Fig. 12, under the challenge of more vague text descriptions and professional dance movements, our dataset can still provide lively hand gestures that align with text controls, demonstrating the broad generalization capabilities of the BOTH57M dataset. Finally, we present a gallery of our inference results for the text/body-to-hand task in Fig. 13.
@@ C Applications
Our BOTH2Hands algorithm, trained on BOTH57M dataset, has demonstrated excellent generalization capabilities even on other human body datasets that only offer body-level descriptions. We conducted experiments by inputting body and text data from the HumanML3D [14] dataset into BOTH2Hands, successfully generating hand movements that are remarkably vivid. In a bid to challenge the generation of hand movements in more professional actions, we proceed to perform inference on InterHuman [29], a dataset encompassing professional movements such as ballet, boxing, and more. Fig. 14 illustrates how our method can effectively augment data for existing body and text datasets.
@@ D Limitation and Broader Impact
In this section, we discuss the border impact and limitations.
Broader Impact. We are the first to propose generating fine-grained two-hand motions through implicit control from real-world body input and user-defined explicit text control. This provides a direction that can be studied for user-customized motion generation. Furthermore, the pre-trained BOTH2Hands, based on the hand synthesis through the parallel diffusion structure and large-scale motion data training on the BOTH57M, can enhance data for large body motion datasets with textual descriptions like HumanML3D [14] and InterHuman [29]. The research also contributes to balancing control signals during the multi-signal control diffusion process. Furthermore, text-controllable and body-aligned hand synthesis has potential in many practical application scenarios, such as Virtual Reality (VR), Augmented Reality (AR) games, animations, and human communication studies, etc. Fig. 15 provides the gallery of data examples captured by our dome with 32 synchronized high-resolution RGB cameras.
Limitation. While this work has achieved substantial advancements in the novel task of generating hand motions in alignment with body and text and offering an intuitive control method, it has limitations. First, the fine-grained spatial control provided by text descriptions is sufficient but lacks temporal alignment methods. Second, we focused on the connection between hand gestures and the body as a whole but did not take into account how different parts of the body might separately affect the hands. Third, more suitable metrics that can precisely reflect the alignment between hand, body, text, and other complex conditions should be presented, we welcome the community to focus on it and hope our contribution will push the field forward. Finally, although BOTH2Hands can generate vivid hand poses, it remains challenging to determine whether a two-handed interaction gesture is generated when the wrists are close enough. Hand-interaction poses represent a hugely different meaning from poses that do not interact. Future work may focus on aligning text control temporally and explore how individual body parts relate to hand movements, especially hand-to-hand interaction. Improved metrics are needed to better represent hand and multi-condition alignment.